### v2

# Sarcasm or Hope???

Abstract
Your abstract.
1 Introduction
Advances in Natural Language Processing (NLP) for emotion recognition have been significant, yet identifying the covert and complex emotion of hope remains a major challenge. Recent improvements in NLP have enhanced applications in sentiment analysis, mental health assessments, social media monitoring, and educational platforms [poria2016deeper, cambria2013new, bustos2021emotion, gomez2023emotion, mishra2025unveiling]. Despite recent progress, a persistent challenge in emotion recognition lies in identifying subtle and complex emotions, particularly hope, which is often overlooked in standard emotion taxonomies [mohammad2021sentiment]. However, these strides have not fully addressed the intricacies of detecting hope. Unlike overt emotions such as anger, fear, sadness or joy, hope is an inherently ambiguous emotion that exhibits differently across various contexts, and carries multiple emotional dimensions [buechel2022emobank, chatterjee2019semeval]. Therefore, traditional NLP techniques struggle to capture subtle emotions and highlight the need for more refined, context-aware approaches to accurately detect nuanced emotional expressions, such as hope [webber2020proceedings].
Accurate detection of hope speech unlocks substantial benefits across real-world NLP applications, such as enhancing student engagement and success [xie2019trends], understanding patients’ mental states [resnik2015beyond], and mental disorder understanding (NPMDU). For example, within online education, instructors can leverage these systems as personalized support tools to accurately identify expressions of hope to better understand students’ motivations, interests, and emotional barriers—ultimately enabling targeted interventions to [xie2019trends]. Similarly, mental health professionals can leverage the natural progression of hope speech to better understand patients’ mental states, therapeutic outcomes, and concealed psychological factors over time. Research emphasizes that accurately interpreting psychological states requires identifying positive emotional expressions, particularly hope and optimism [resnik2015beyond]. As a result, NLP systems significantly benefit from incorporating hope detection technologies, since integrating positive emotional indicators alongside negative sentiments enables more comprehensive, nuanced, and effective analytic models.
Traditional NLP techniques struggle to detect hope due to its complex, dual nature as defined by Lazarus [lazarus1991emotion]. According to Lazarus, hope consists of two essential components: (i) a specific desire for positive outcomes, and (ii) an overall optimistic view of future events. The inherent contradictions within hope result in complicated textual interpretations, causing individual textual markers to be variably perceived as genuine goals, imaginary objectives, or sardonic remarks. Consequently, expressions intended to reflect hopeful thinking may inadvertently suggest agitation, depending on their textual context. For example, the sentence “I hope things finally change around here” reveals optimism but also hints at frustration, depending on how it is used in the text. Most emotion recognition systems trained on simplified affective categories (e.g., Ekman’s six basic emotions) tend to fail when processing hope, as a generic positive sentiment or as non-committal discourse, failing to grasp its subtleties [ekman1992argument, buechel2022emobank]. Thus, emotion recognition systems trained on basic emotion sets fail to correctly identify hope and often misclassify it either as a simple positive emotion or as a neutral, non-committal state.
Additionally, recognizing hope is further complicated by sarcasm, as sarcastic expressions frequently employ positive language to communicate negative sentiments. The mismatch between literal meanings and intended emotions makes it challenging for machine learning and deep learning classifiers to distinguish genuinely hopeful statements from their sarcastic counterparts. For instance, sarcasm occurs when someone says, “Great, I really hope this mess gets worse,” even though the words themselves convey hope, but the semantic context reveals sarcastic negativity [riloff2013sarcasm].
The semantic incongruity between lexical content and emotional meaning makes emotion detection more challenging. The combination of figurative language and sarcasm complicates consistent emotion classification for computers since they primarily work with superficial word analysis instead of contextual interpretation [poria2016deeper]. Accurate context-aware emotion detection primarily depends on extensive, diverse datasets and sophisticated models [cambria2017sentiment].
However, there is a notable lack of annotated datasets explicitly related to hope, and existing datasets often either entirely omit hope or classify it merely as a subset of general positive emotions that cause incorrect predictions. Therefore, more comprehensive annotated datasets are necessary that explicitly capture hope through genuinely positive sentiments, unfounded expectations, and sarcastic expressions.
In this study, we bridge the gap between emotion recognition and pragmatic language understanding and present a bilingual annotated hope speech dataset designed for hope detection in both English and Spanish language. We explicitly annotate sarcasm within hope expressions and create rich annotation guidelines for different hope categories. We design four new categories (i) generalized optimism, (ii) realistic expectations, (iii) unrealistic or wishful thinking, and (iv) sarcastic hope. To ensure accuracy and consistency in annotations, we rigorously adhered to detailed taxonomical guidelines and employed inter-annotator agreement protocols. We develop context-aware, sarcasm-resilient emotion classification models as a baseline to detect and evaluate the performance of the hope detection system in English and Spanish languages. This dataset can be used for advancing cross-linguistic emotion recognition research.
2 Literature Review
Natural Language Processing (NLP) has advanced beyond basic sentiment classification (positive, negative, or neutral) toward identifying more nuanced emotional states, including the subtle and underexplored emotion of hope [cambria2017sentiment, mohammad2021sentiment]. While emotions such as joy, fear, and regret have received substantial attention, hope remains comparatively underresearched, despite its significant role in mental health support [resnik2015beyond], education [xie2019trends], online communication, and social programs [ullah2024hope]. Hope consists of two distinct components that contribute to its complexity: (i) a general optimism inherent in individuals, and (ii) a specific expectation for positive outcomes [lazarus1991emotion]. Given its unique semantic, emotional, and contextual characteristics, recent research in hope speech detection has increasingly treated hope as an independent emotional category rather than merely grouping it under general positive emotions.
2.1 Methodologies for Hope Speech Detection
Different methodologies have been proposed for detecting hope speech and addressing challenges related to context, semantics, and multilingual content. Various studies have explored techniques to identify hope expressions within social media and textual documents [strapparava2008learning, palakodety2020hope]. For instance, Balouchzahi et al. [balouchzahi2023polyhope] introduced “PolyHope,” an NLP system that operates at semantic and contextual levels to detect nuanced expressions of hope. Similarly, another study [sidorov2023regret] leveraged transformer-based models to differentiate between expressions of hope and regret, and highlighted complexities in jointly modeling these emotionally similar pairs. Despite these advancements, current methodologies lack in distinguishing hope expressions from sarcastic or superficially hopeful messages, particularly on social media platforms (e.g., X (previously Twitter), and Reddit). Deep learning models provide poor results because they often misinterpret pragmatic signals and contextual discourse markers [taboada2016sentiment]. These limitations underline the need to incorporate pragmatic and contextual awareness to improve model accuracy. Moreover, addressing these contextual challenges becomes even more critical when extending hope detection methodologies across multilingual and cross-cultural datasets.
A growing body of work has explored multilingual and cross-lingual hope speech detection, recognizing its importance in linguistically diverse societies. Chakravarthi et al. [chakravarthi2020hopeedi, chakravarthi2022overview] introduced multilingual approaches to detect hope speech in English, Tamil, Malayalam, and Kannada. The authors explicitly emphasized equality and diversity and inclusion (EDI) to detect hope speech. Nevertheless, their models did not sufficiently account for distortions caused by sarcasm or unrealistic expectations embedded within hopeful expressions. Similarly, researchers such as Arunima et al. [arunima2021ssn_dibertsity] have contributed to shared tasks like Hope Speech Detection for Equality, Diversity, and Inclusion (HopeEDI), which expanded datasets and evaluation benchmarks in multiple Indian languages. Additional efforts have targeted Urdu datasets, with studies exploring the transferability of hope speech models across languages and the need for language-specific semantic resources [balouchzahi2025urduhope, ponnusamy2024vel]. Chakravarti [chakravarthi2022multilingual] addressed hope speech detection in Tamil, English, and Malayalam, highlighting classification challenges due to code-mixed data. Moreover, Malik et al. [malik2023multilingual] explored a joint multilingual and translation-based approach for hope speech detection in English and Russian using a fine-tuned Russian-RoBERTa model, achieving 94% accuracy and 80.24However, most of these models do not explicitly address the confounding effects of sarcasm, exaggeration, or unrealistic optimism expressed through hopeful language—factors that often distort detection accuracy in real-world applications. Moreover, research on low-resource and code-mixed language scenarios remains limited, although recent work has begun to tackle these challenges through cross-lingual transfer learning and zero-shot classification approaches [thangaraj2024cross].
Moreover, the integration of sarcasm detection and contextual awareness into current lexical and dictionary-based hope detection methods often underperformed in analyzing unstructured, informal textual data. Therefore, integrating pragmatic, contextual, and multilingual perspectives to ensure accurate and reliable emotion recognition outcomes is crucial.
2.2 Existing Datasets for Hope Speech Detection
Several datasets have been introduced to facilitate advancements in hope speech detection in various languages, such as Spanish [garcia2023hope, garcia2024overview], Bengali [nath2023bonghope]. For example, the “HopeEDI” dataset [chakravarthi2020hopeedi] introduced a multilingual corpus specifically targeting hopeful expressions related to equality, diversity, and inclusion (EDI). Additionally, the “IberLEF” Hope datasets [garcia2023hope, garcia2024overview] extended hope speech detection into Spanish, and emphasized annotations around social causes and community identities. Another dataset, “BongHope” [nath2023bonghope], presented a Bengali corpus for hope speech detection. Furthermore, the HOPE shared tasks at IberLEF provided comprehensive benchmarks and detailed insights into effective methodologies for annotating hope datasets [garcia2024overview, jimenez2023overview]. In parallel, the HASOC-Hope subtrack, part of the FIRE shared tasks, introduced additional multilingual datasets covering Hindi, Malayalam, and Tamil, further enriching the research landscape in this area [mandl2020overview]). Finally, the LREC Hope Speech dataset offers a multilingual resource tailored for cross-lingual and zero-shot hope speech detection, with a special focus on low-resource language scenarios [pelicon2021zero].
However, most existing datasets either categorize hope speech through a simplistic binary labeling scheme (hope vs. non-hope) or do not annotate nuanced subcategories such as generalized optimism, unrealistic hope, or sarcastic hope. Additionally, few datasets explicitly annotate sarcasm, which limits their usefulness in training or evaluating models capable of differentiating genuine hope from sarcastic expressions. Moreover, annotation schemas typically lack cross-linguistic alignment, which poses challenges for developing robust multilingual models. As a result, there remains a notable gap in available datasets that provide both fine-grained hope annotations and explicit sarcasm labeling.
2.3 Sarcasm in Hope Speech: A Complex Challenge
The detection of hope speech is highly impeded by the presence of sarcasm. Positive sentimental language through rhetorical expression functions as sarcasm to convey negative meanings while possessing a humorous or mocking quality. For example, “This one needs to go well as it happened previously.” This hopeful-sounding statement contains semantic skepticism through ironic speech. The model detection of hopeful speech can fail due to sarcastic statements that resemble hopeful statements. The improper evaluation of sentiment features occurs when basic analysis methods, such as positive polarity measurements, operate independently from detailed contextual analysis.
The task of sarcasm detection is challenging because it requires contextual information, tonal cues, and background understanding [cambria2017sentiment]. Research on sarcasm detection within NLP has approached this problem in four ways: (i) Contextual embeddings (e.g., BERT, RoBERTa), (ii) Multitask learning frameworks (to jointly learn emotion and sarcasm), (iii) Contrastive learning approaches, and (iv) Discourse-aware models that incorporate surrounding sentences or conversational history. Yet sarcasm-aware hope detection remains an unexplored intersection, and no existing hope detection dataset provides labeled sarcastic examples, as emphasized by Balouchzahi et al. [balouchzahi2023polyhope] in their analysis of NPMDU-related textual data. To address this challenge, we annotate hope expressions along with sarcasm indicators and provide a richer, more realistic representation of hope in social discourse. Such annotation is critical for building models that can distinguish between sincere emotional optimism and sarcastic cynicism, as such distinction is particularly important in domains like education, counseling, and mental health.
3 Methodology
This section details the baselines employed for hope speech and sarcasm detection within the PolyHope v.2 dataset. We explore two primary approaches: fine-tuning pre-trained Transformer models and leveraging Zero-Shot Learning (ZSL) and Few-Shot Learning (FSL) capabilities of Large Language Models (LLMs).
3.1 Fine-tuning Pre-trained Transformer Models
To establish strong and reproducible baselines for the proposed task, we adopted a fine-tuning approach using pre-trained Transformer models from the Hugging Face Transformers library  [^1] https://huggingface.co/models. This approach leverages transfer learning, enabling us to adapt models pre-trained on large corpora to our task-specific dataset. We fine-tuned several Transformer architectures, including Roberta, Albert, Electra, and DistilBERT, to evaluate their performance on both binary (Hope vs. Not Hope) and multiclass (Not Hope, Generalized Hope, Realistic Hope, Unrealistic Hope, and Sarcasm) classification tasks.
3.1.1 Experimental Setup
We implemented a stratified 5-fold cross-validation strategy using the simpletransformers library  [^2] http://simpletransformers.ai/ to ensure robust and generalizable evaluation. This is crucial to mitigate potential biases and provide reliable performance estimates. Within each fold, we executed the following steps:
  1. 1.
    Model Initialization: We instantiated a ClassificationModel from the simpletransformers library, specifying the pre-trained architecture and model variant.
  2. 2.
    Fine-tuning Procedure: The instantiated model was then fine-tuned on the training partition. We employed early stopping based on the development set performance to prevent overfitting.
  3. 3.
    Evaluation Metrics: After fine-tuning, the model’s performance was assessed on the held-out test split. We report accuracy, weighted precision, weighted recall, weighted F1-score, macro precision, macro recall, and macro F1-score. We emphasize weighted and macro averages due to the imbalanced nature of the dataset, providing a comprehensive performance profile across all classes.
  4. 4.
    Qualitative Analysis: To gain deeper insights into model behavior, we generated classification reports and confusion matrices for each fold. The confusion matrices were visualized to enable identification of systematic misclassifications. These matrices were averaged across folds to provide an aggregate view of model confusions.
3.1.2 Model Selection and Training Regimen
We conducted experiments on both English and Spanish datasets. The specific models used for each language are detailed in Tables 1 and 2. All models were fine-tuned using the AdamW optimizer with a learning rate of 3e-5 and a weight decay of 0.01. We trained for a maximum of 15 epochs per fold, employing early stopping with a patience of 3 epochs based on the performance on a held-out development set. The maximum sequence length was set to 100 tokens. A batch size of 32 was used for all experiments. The choice of these hyperparameters was guided by a coarse grid search on a separate development set, aiming to optimize performance while preventing overfitting.
Table 1: Pre-trained Transformer Models Used for English
Architecture Model Name
Roberta roberta-base
Albert albert-base-v2
Electra google/electra-base-generator
DistilBERT distilbert-base-uncased
Table 2: Pre-trained Transformer Models Used for Spanish
Architecture Model Name
Roberta PlanTL-GOB-ES/roberta-base-bne
Albert dccuchile/albert-base-spanish
Electra mrm8488/electricidad-base-generator
DistilBERT mrm8488/distill-bert-base-spanish-wwm-cased-finetuned-spa-squad2-es
Through this fine-tuning exercise, we aim to establish strong, reproducible baselines for the PolyHope task. The results obtained will quantify the effectiveness of Transformer models for hope speech detection and, critically, sarcasm identification. This comparative analysis will provide a benchmark against which to evaluate the performance gains achieved by our proposed ZSL and FSL methodologies leveraging large language models.
3.2 Zero-Shot and Few-Shot Learning with Large Language Models
We further investigated the potential of LLMs to perform hope speech and sarcasm detection in ZSL and FSL settings. This approach allows us to leverage the pre-existing knowledge and reasoning capabilities of LLMs without extensive task-specific training. We employed GPT-4 and Llama 3, two state-of-the-art LLMs, to explore their effectiveness in this context.
We selected GPT-4 [achiam2023gpt, bubeck2023sparks] and Llama 3 [grattafiori2024llama] for their superior performance in various natural language understanding and generation tasks. GPT-4, known for its advanced reasoning and generalization abilities, serves as a high-performing benchmark. Llama 3, a powerful open-source LLM, offers an alternative perspective and allows for greater control and customization.
3.2.1 Prompt Engineering
The performance of LLMs in ZSL and FSL settings is highly dependent on the design of effective prompts. We carefully crafted prompts to elicit the desired classification behavior from the LLMs.
Zero-Shot Prompts: For ZSL, we used a simple prompt that instructed the LLM to classify the input text into one of the predefined categories (Hope, Not Hope for binary classification; Generalized Hope, Realistic Hope, Unrealistic Hope, Not Hope, Sarcasm for multiclass classification). The prompt was formatted as follows:
Classify the following text into one of the $\{categories\}$:
[…]
Label:
Few-Shot Prompts: For FSL, we provided the LLMs with a set of labeled examples to guide their classification. We selected 5 balanced examples per class. The prompt was structured as follows:
Here are some labeled examples:
[…]
Now classify this text:
[…]
Label:
3.2.2 Implementation Details
GPT-4: We used the OpenAI API to access GPT-4. The classify_text function sends a request to the OpenAI API with the appropriate prompt and receives the LLM’s classification. The temperature parameter was set to 0 to encourage deterministic and consistent responses.
response = openai.chat.completions.create(
    model="gpt-4",
    messages=[{"role": "system", "content": "You are an NLP model trained for text classification."},
              {"role": "user", "content": prompt}],
    temperature=0
Llama 3: We accessed Llama 3 via the Together AI API. The format_few_shot_prompt function constructs the prompt, including a system message defining the LLM’s role as a text classification expert and providing example classifications. The Together AI API client then sends the prompt to the Llama 3 model. We set the temperature to 0.3 and the maximum number of tokens to 50.
response = client.chat.completions.create(
    model="meta-llama/Llama-3-70b-chat-hf",
    messages=messages,
    temperature=0.3,
    max_tokens=50,
[…]
    stream=True
To evaluate the performance of both GPT-4 and Llama 3, we used the same evaluation metrics as in the fine-tuning experiments. These metrics allow for a direct comparison between the LLM-based approaches and the fine-tuned Transformer baselines.
By evaluating GPT-4 and Llama 3 in ZSL and FSL settings, we aim to assess the potential of LLMs for hope speech and sarcasm detection without task-specific training. The results will provide insights into the strengths and limitations of these models and inform future research directions in this area. We expect that FSL will outperform ZSL, and that GPT-4 and Llama 3 will exhibit different strengths and weaknesses due to their architectural differences and training data. A comparison with the fine-tuned Transformer baselines will reveal the trade-offs between these two approaches and highlight the potential for combining them to achieve even better performance.